# A hybrid electromechanical phase-field and deep learning framework for predicting fracture in dielectric nanocomposites

### v1

@@ 3.2 Training and validation
The networks were trained separately on two data modalities derived from physics-based FE simulations: (i) phase-field scalar field ($\mathfrak{d}$) and (ii) electric potential ($\phi$). Each dataset comprised 10,000 unique grayscale images of size $572\times 572$ pixels, representing 2D slices of simulated specimens with randomly distributed defects. The crack paths captured via phase-field or electric potential gradients served as ground-truth labels, enabling supervised segmentation learning.
All networks were implemented using TensorFlow and trained on an NVIDIA RTX 3090 GPU with 24 GB of memory. The training process employed the Adam optimizer with an initial learning rate of $1\times 10^{-4}$, using $\beta_{1}=0.9$ and $\beta_{2}=0.999$ to control the decay rates for the moment estimates. An adaptive learning rate scheduler was used to reduce the learning rate by a factor of 0.1 if the validation loss plateaued. To prevent overfitting, early stopping was applied. The loss function was a composite of 50% Dice loss and 50% binary cross-entropy loss, allowing the model to balance region-level accuracy with pixel-wise classification. Training was conducted over a maximum of 100 epochs, subject to early stopping criteria, using a batch size of 24. To enhance generalization, data augmentation was applied in the form of random rotations ($\pm 15^{\circ}$), horizontal and vertical flips, and additive Gaussian noise with standard deviation $\sigma=0.05$, see [46, 47]. The dataset was randomly partitioned into 80% training, 10% validation, and 10% testing subsets using stratified sampling to maintain label distribution.
As a representative example, the training and validation metrics for the ResNet34-based U-Net model are shown in Figure 4 for the for phase-field data. Initially, the model yielded an F1-score of 0.75 and an IoU of 0.67, with a training loss of 0.27. Over the course of 75 epochs, the model achieved an F1-score of 0.89, IoU of 0.82, and a reduced training loss of 0.12, with corresponding validation metrics stabilizing at 0.82, 0.75, and 0.19 respectively.
@@
The consistent reduction in loss and concurrent increase in F1 and IoU scores for both data types confirms that the networks were effectively learning spatial correlations and developing strong crack segmentation capabilities. Importantly, the models trained on electric potential images exhibited faster convergence and superior validation metrics, likely due to the smoother and more continuous field structure in electric potential distributions, which may offer richer gradient information to CNNs compared to localized phase-field inputs.
@@ 3.3 Performance with phase-feld data
Herein, the CNN models were trained to learn and segment crack paths directly from input phase-field images using four ResNet backbones: ResNet34, ResNet50, ResNet101, and ResNet152. Each architecture was evaluated over a wide range of simulation cases with varying defect distributions and crack topologies.
Table 2 summarizes the performance of the models on the phase-field dataset. Among the architectures, ResNet34 provided the best trade-off between accuracy and computational efficiency. This architecture reached a training F1-score of 0.89 and an IoU of 0.82 after 75 epochs, with a total training time of approximately 2 hours. On the validation set, ResNet34 reached an F1-score of 0.82 and an IoU of 0.75, with a final validation loss of 0.19.
@@
Table 2: CNN Performance on phase-field input data.
@@
The improvement in IoU over epochs is visualized in Figure 4 (middle), while the training/validation loss curve in Figure 4 (left) confirms a stable convergence pattern. Notably, the phase-field models required a larger number of training iterations to converge compared to the electrical potential models, indicating the presence of sharper gradients and higher sensitivity in crack tip regions.
Figures 6 and 7 illustrate two cases–Case 1 and Case 2, respectively–demonstrating the sequential process of phase-field fracture detection and labeling. Each figure presents a series of three images. The original image (left) serves as the baseline input for analysis. The phase-field crack image (middle) shows the application of phase-field modeling to identify crack patterns within the material domain. The final image (right), referred to as the label (mask) image, displays a binary mask that clearly delineates the exact locations of the detected cracks. This sequence underscores the importance of each step in the detection pipeline and effectively visualizes how phase-field techniques are employed for crack identification and annotation.
Figure 6: Case 1 – (Left) Original image, (Middle) Phase-field crack image, and (Right) Label image.
Figure 7: Case 2 – (Left) Original image, (Middle) Phase-field crack image, and (Right) Label image.
@@
As an example, in Case 1, the predicted crack pattern closely aligns with the ground truth label, particularly in regions with branching behavior. This indicates the model’s effectiveness in capturing complex crack propagation paths. However, some blurring was observed at the crack tips, which is likely attributed to limitations in the receptive field resolution of the convolutional network. The confusion matrices presented in Figures 8 through 15 further support this observation, showing relatively low false positive rates across models, but highlighting the presence of false negatives in micro-cracked regions, especially where cracks are thinner and more dispersed.
@@
3.4 Performance with electrical potential data
Electrical potential data arise from the electrostatic field solution within the electromechanical phase-field framework. In dielectric materials, crack formation disturbs the continuity of the electric field, manifesting as gradients and distortions in the potential map. Unlike the highly localized phase-field field, electric potential images are smoother and contain distributed features, providing implicit cues about crack presence through field variation rather than direct damage encoding.
@@ 3.4 Performance with electrical potential data
Figure 5 shows the loss and IoU trends during training, revealing a faster and more stable convergence than phase-field-based models. The electrical potential-based models reached near-plateau performance in under 40 epochs, reflecting the information richness of the electric potential field for guiding crack prediction.
Table 3: CNN Performance on electrical potential input data.
@@
Figure 16: Case 1 – (Left) Original image, (Middle) Electrical potential crack image, and (Right) Label image.
Figure 17: Case 2 – (Left) Original image, (Middle) Electrical potential crack image, and (Right) Label image.
Figure 18: Case 1 – (Left) Predicted crack image for Resnet34 based on electrical potential data and (Right) Confusion matrix.
Figure 19: Case 2 – (Left) Predicted crack image for Resnet34 based on electrical potential data and (Right) Confusion matrix.
Figure 20: Case 1 – (Left) Predicted crack image for ResNet50 based on electrical potential data and (Right) Confusion matrix.
Figure 21: Case 2 – (Left) Predicted crack image for ResNet50 based on electrical potential data and (Right) Confusion matrix.
Figure 22: Case 1 – (Left) Predicted crack image for ResNet101 based on electrical potentialdata and (Right) Confusion matrix.
Figure 23: Case 2 – (Left) Predicted crack image for ResNet101 based on electrical potential data and (Right) Confusion matrix.
Figure 24: Case 1 – (Left) Predicted crack image for ResNet152 based on electrical potential data and (Right) Confusion matrix.
Figure 25: Case 2 – (Left) Predicted crack image for ResNet152 based on electrical potential data and (Right) Confusion matrix.
Similarly, Figures 16 and 17 present the image preprocessing pipeline for two representative cases–Case 1 and Case 2–used to generate ground truth crack labels from electric potential simulations. Each figure illustrates three key stages in the data preparation process. The first image (left) shows the original image, which captures the distribution of holes. The second image (middle) shows the corresponding electric potential map, which reveals spatial voltage gradients induced by the presence of cracks. The final image (right) is the binary mask label, where regions of crack-induced electric field distortion are converted into clear segmentation boundaries. This sequential structure effectively highlights the transformation from continuous voltage fields into discrete segmentation labels, which the CNN learns to replicate.
Figures 18 through 25 demonstrate the predictive performance of ResNet-based U-Net architectures for both Case 1 and Case 2 when trained on electrical potential data. Each figure consists of two subcomponents: the left panel shows the predicted crack region based on CNN inference, while the right panel presents the corresponding confusion matrix. Specifically, Figures 18 and 19 showcase ResNet34 results, Figures 20 and 21 correspond to ResNet50, Figures 22 and 23 illustrate ResNet101, and Figures 24 and 25 present the outputs from ResNet152. These visualizations provide clear comparative insights into the architectural impacts on crack localization accuracy, generalization, and robustness across input conditions.
@@
Electric potential images offer several advantages as training inputs. First, their inherently smooth and continuous nature helps convolutional filters extract long-range spatial features without being misled by pixel-level noise. Second, distortions in the voltage gradient field, while subtle, are often more spatially extensive than discrete crack fronts, providing CNNs with a broader contextual field to base decisions upon. Third, because electric potential perturbations result from the disruption of conductive pathways due to cracking, they implicitly encode crack severity and propagation direction. This richness in physical interpretation makes them particularly suited for deep learning-based surrogate modeling.
@@ 3.5 Comparative insight
The preceding sections demonstrated the predictive capabilities of deep convolutional networks trained on two distinct types of input modalities: phase-field fracture fields and electrical potential fields. While both datasets originate from the same underlying finite element simulations of dielectric nanocomposites subjected to mechanical and electrical loading, they represent fundamentally different physical observables. The phase-field field directly encodes the material’s internal damage state as a scalar damage variable, whereas the electrical potential field reflects the system’s electrostatic response, perturbed indirectly by the presence and evolution of cracks. This section provides a comparative evaluation of the two approaches, grounded in performance metrics, architectural sensitivity, interpretability, and practical implications for real-world applications.
@@
Tables 2 and 3 summarize the training and validation performance of models ranging from ResNet34 to ResNet152, using both phase-field data and electrical potential inputs, respectively. As evident from these tables, models trained on electrical potential data consistently outperformed their phase-field-based counterparts across all ResNet backbones. They not only achieved higher validation F1 and IoU scores but also demonstrated faster and more stable convergence. For instance, the ResNet34 model trained on electrical potential data reached a validation IoU of 0.89 within just 50 epochs, whereas the ResNet34 model trained on phase-field data plateaued at an IoU of 0.76 after 75 epochs. This suggests that the electrical potential fields provided richer, more learnable features for crack segmentation.
@@
A key insight from the comparative analysis is the role of network depth in segmentation performance across different input modalities. When trained on electrical potential data, deeper ResNet variants–such as ResNet101 and ResNet152–show marginal but consistent improvements in validation accuracy compared to ResNet34. For instance, ResNet152 achieved a validation F1-score of 0.94 and a validation IoU of 0.90, slightly outperforming ResNet34, which achieved 0.93 and 0.89, respectively. These results indicate that deeper architectures can extract and utilize subtle spatial cues embedded in the smooth gradients of electrical potential fields.
@@
In contrast, no such improvement is observed with deeper networks when trained on phase-field data. Despite the higher representational capacity of ResNet101 and ResNet152, their performance on phase-field inputs either stagnates or slightly declines relative to ResNet34. For example, ResNet34 outperformed all other variants with a validation F1-score of 0.82 and an IoU of 0.75, while deeper models failed to surpass this benchmark despite significantly longer training times. This suggests that the highly localized and discontinuous nature of phase-field crack images may not benefit from deeper architectures, which are generally more effective at learning hierarchical or global representations.
@@
The implication is that for phase-field-based segmentation tasks, increasing model depth yields diminishing or even negative returns, potentially due to overfitting to noise or sharp transitions in the data. On the other hand, electrical potential data–with their smooth, continuous gradients–support better feature extraction in deeper networks, albeit with marginal gains.
Interpretability and physics-informed learning
@@
While the phase-field approach is more directly rooted in fracture modeling from a mechanics perspective, its effectiveness in training convolutional neural networks (CNNs) is somewhat constrained by its highly localized nature. In contrast, electrical potential data represent a smoothed, global field that is influenced not only by material fracture but also by the underlying geometry and boundary conditions. This yields several advantages for deep learning. Electrical potential fields offer context-rich features by providing diffuse gradients that inherently reflect the influence of cracks, thereby enabling the network to better learn spatial dependencies. Moreover, because electric field distortions extend beyond the immediate vicinity of the crack, the data provide greater spatial coverage, improving class balance and leading to more stable CNN training. Importantly, electrical potential fields naturally embed electro-mechanical coupling effects, allowing the network to learn more comprehensive representations that capture not only where damage occurs but also how it affects the overall system behavior.
@@
Despite its superior performance, electrical potential data-based learning is not without limitations. One challenge lies in interpretability: while phase-field values directly correlate with damage, interpreting voltage-induced field distortions as ”cracks” requires physical insight and a well-trained model. Furthermore, in complex geometries or low signal-to-noise regimes, voltage gradients may be too diffuse to pinpoint precise crack boundaries, potentially affecting precision.

### v2

@@ 3.2 Training and validation
The networks were trained separately on two data modalities derived from physics-based FE simulations: (i) the phase-field scalar field ($\mathfrak{d}$) and (ii) the electric potential ($\phi$). Each dataset comprised 10,000 unique single-channel images of size $572\times 572$ pixels, representing 2D slices of simulated specimens with randomly distributed defects. Since the finite element simulations utilize an unstructured mesh to capture the electromechanical response, a post-processing rasterization step was performed to interpolate the nodal field data onto this structured Cartesian grid to ensure compatibility with the CNN architecture.
A critical distinction between these modalities is their numerical range; while the phase-field variable $\mathfrak{d}$ is inherently bounded in $[0,1]$, the raw electric potential $\phi$ is physically dependent on the applied boundary conditions and material properties. To ensure architectural consistency and training stability, both modalities were normalized to a common range of $[0,1]$ before being passed to the ResNet-U-Net. Specifically, the electric potential fields were linearly rescaled for each sample such that the local minimum and maximum values map to 0 and 1, respectively. This min-max normalization makes the network invariant to absolute voltage levels and ensures that the training statistics are consistent across all training samples. While finite element results are often visualized as RGB contour plots for illustrative purposes (see Figure 3), the actual data used for training and inference are strictly normalized grayscale images to optimize computational efficiency. The crack paths captured via phase-field or electric potential gradients served as the ground-truth labels for supervised segmentation learning.
All networks were implemented using TensorFlow and trained on an NVIDIA RTX 3090 GPU with 24 GB of memory. The training process employed the Adam optimizer with an initial learning rate of $1\times 10^{-4}$, using $\beta_{1}=0.9$ and $\beta_{2}=0.999$ to control the decay rates for the moment estimates. An adaptive learning rate scheduler was used to reduce the learning rate by a factor of 0.1 if the validation loss plateaued. To prevent overfitting, early stopping was applied. The loss function was a composite of 50% Dice loss and 50% binary cross-entropy loss, allowing the model to balance region-level accuracy with pixel-wise classification. Training was conducted over a maximum of 100 epochs, subject to early stopping criteria, using a batch size of 24. To enhance generalization, data augmentation was applied in the form of random rotations ($\pm 15^{\circ}$), horizontal and vertical flips, and additive Gaussian noise with standard deviation $\sigma=0.05$, see Mikołajczyk and Grochowski [2018], Mumuni and Mumuni [2022]. The dataset was randomly partitioned into 80% training, 10% validation, and 10% testing subsets using stratified sampling to maintain label distribution.
As a representative example, the training and validation metrics for the ResNet34-based U-Net model are shown in Figure 4 for the phase-field data. Initially, the model yielded an F1-score of 0.75 and an IoU of 0.67, with a training loss of 0.27. Over the course of 75 epochs, the model achieved an F1-score of 0.89, IoU of 0.82, and a reduced training loss of 0.12, with corresponding validation metrics stabilizing at 0.82, 0.75, and 0.19 respectively.
@@
The consistent reduction in loss and concurrent increase in F1 and IoU scores for both data types confirms that the networks were effectively learning spatial correlations and developing strong crack segmentation capabilities. Importantly, the model trained using electric potential signatures exhibited faster convergence and superior validation metrics, likely due to the smoother and more continuous field structure in electric potential distributions, which may offer richer gradient information to CNNs compared to localized phase-field data.
@@ 3.3 Performance with phase-feld data
Herein, the CNN models were trained to learn and segment crack paths directly from phase-field data using four ResNet backbones: ResNet34, ResNet50, ResNet101, and ResNet152. Each architecture was evaluated over a wide range of simulation cases with varying defect distributions and crack topologies.
Table 3 summarizes the performance of the models on the phase-field dataset. The performance metrics in this table are derived from a pixel-wise confusion matrix analysis, comparing every pixel of the CNN prediction against the ground-truth data. Among the architectures, ResNet34 provided the best trade-off between accuracy and computational efficiency. This architecture reached a training F1-score of 0.89 and an IoU of 0.82 after 75 epochs, with a total training time of approximately 2 hours. On the validation set, ResNet34 reached an F1-score of 0.82 and an IoU of 0.75, with a final validation loss of 0.19.
@@
Table 3: CNN Performance on phase-field data.
@@
The improvement in IoU over epochs is visualized in Figure 4 (middle), while the training/validation loss curve in Figure 4 (left) confirms a stable convergence pattern. Notably, the phase-field models required a larger number of epochs to converge compared to the electric potential models, indicating the presence of sharper gradients and higher sensitivity in crack tip regions.
Figures 6 and 7 illustrate two cases–Case 1 and Case 2, respectively–demonstrating the the relationship between the physical simulation and the training data. The left panels show the original input images provided to the CNN, which capture the specific geometric distribution of randomly placed circular holes. The middle panels display the actual crack paths (ground-truth) derived directly from the phase-field damage variable ($\mathfrak{d}$), while the right-most panels show the binary crack masks used as labels for CNN training.
Figure 6: Case 1 – Image preprocessing pipeline. (Left) Original input image showing the random distribution of holes, (Middle) Ground-truth crack path obtained from the phase-field damage variable in high-fidelity FE simulations, and (Right) Corresponding binary crack masks used as labels for training the deep learning model.
Figure 7: Case 2 – Image preprocessing pipeline. (Left) Original input image showing the random distribution of holes, (Middle) Ground-truth crack path obtained from the phase-field damage variable in high-fidelity FE simulations, and (Right) Corresponding binary crack masks used as labels for training the deep learning model.
@@
As an example, in Case 1, the predicted crack pattern closely aligns with the ground truth label, particularly in regions with branching behavior. This indicates the model’s effectiveness in capturing complex crack propagation paths. However, some blurring was observed at the crack tips, which is likely attributed to limitations in the receptive field resolution of the convolutional neural network. The confusion matrices presented in Figures 8 through 15 further support this observation, showing relatively low false positive rates across models, but highlighting the presence of false negatives in micro-cracked regions, especially where cracks are thinner and more dispersed.
@@
3.4 Performance with electric potential data
Electric potential data arise from the electrostatic field solution within the electromechanical phase-field framework. In dielectric materials, crack formation disturbs the continuity of the electric field, manifesting as gradients and distortions in the potential map. Unlike the highly localized phase-field field, electric potential images are smoother and contain distributed features, providing implicit cues about crack presence through field variation rather than direct damage encoding.
@@ 3.4 Performance with electric potential data
Figure 5 shows the loss and IoU trends during training, revealing a faster and more stable convergence than phase-field-based models. The electric potential-based models reached near-plateau performance in under 40 epochs, reflecting the information richness of the electric potential field for guiding crack prediction.
Table 4: CNN Performance on electric potential data.
@@
Figure 16: Case 1 – Image preprocessing pipeline. (Left) Original input image showing the random distribution of holes, (Middle) Ground-truth crack path obtained from the electric potential field in high-fidelity FE simulations, and (Right) Corresponding binary crack masks used as labels for training the deep learning model.
Figure 17: Case 2 – Image preprocessing pipeline. (Left) Original input image showing the random distribution of holes, (Middle) Ground-truth crack path obtained from the electric potential field in high-fidelity FE simulations, and (Right) Corresponding binary crack masks used as labels for training the deep learning model.
Figure 18: Case 1 – (Left) Predicted crack image for Resnet34 based on electric potential data and (Right) Confusion matrix.
Figure 19: Case 2 – (Left) Predicted crack image for Resnet34 based on electric potential data and (Right) Confusion matrix.
Figure 20: Case 1 – (Left) Predicted crack image for ResNet50 based on electric potential data and (Right) Confusion matrix.
Figure 21: Case 2 – (Left) Predicted crack image for ResNet50 based on electric potential data and (Right) Confusion matrix.
Figure 22: Case 1 – (Left) Predicted crack image for ResNet101 based on electric potentialdata and (Right) Confusion matrix.
Figure 23: Case 2 – (Left) Predicted crack image for ResNet101 based on electric potential data and (Right) Confusion matrix.
Figure 24: Case 1 – (Left) Predicted crack image for ResNet152 based on electric potential data and (Right) Confusion matrix.
Figure 25: Case 2 – (Left) Predicted crack image for ResNet152 based on electric potential data and (Right) Confusion matrix.
Similarly, Figures 16 and 17 present the image preprocessing pipeline for two representative cases–Case 1 and Case 2–used to generate ground truth crack labels from electric potential simulations. Each figure illustrates three key stages in the data preparation process. The first image (left) shows the original input image, which captures the distribution of holes. The second image (middle) shows the corresponding high-fidelity electric potential map, which reveals spatial voltage gradients induced by the presence of cracks. The final image (right) is the binary mask label, where regions of crack-induced electric field distortion are converted into clear segmentation boundaries.
Figures 18 through 25 demonstrate the predictive performance of ResNet-based U-Net architectures for both Case 1 and Case 2 when trained on electric potential data. Each figure consists of two subcomponents: the left panel shows the predicted crack region based on CNN inference, while the right panel presents the corresponding confusion matrix. Specifically, Figures 18 and 19 showcase ResNet34 results, Figures 20 and 21 correspond to ResNet50, Figures 22 and 23 illustrate ResNet101, and Figures 24 and 25 present the outputs from ResNet152. These visualizations provide clear comparative insights into the architectural impacts on crack localization accuracy, generalization, and robustness across input conditions.
@@
Utilizing electric potential fields as physical signatures during training offers several advantages. First, their inherently smooth and continuous nature helps convolutional filters extract long-range spatial features without being misled by pixel-level noise. Second, distortions in the voltage gradient field, while subtle, are often more spatially extensive than discrete crack fronts, providing CNNs with a broader contextual field to base decisions upon. Third, because electric potential perturbations result from the disruption of conductive pathways due to cracking, they implicitly encode crack severity and propagation direction. This richness in physical interpretation makes them particularly suited for deep learning-based surrogate modeling.
@@ 3.5 Comparative insight
The preceding sections demonstrated the predictive capabilities of deep convolutional neural networks trained using two distinct physical signatures to guide the mapping from geometry to fracture: phase-field damage variables and electric potential fields. While both datasets originate from the same underlying finite element simulations of dielectric nanocomposites subjected to mechanical and electrical loading, they represent fundamentally different physical observables. The phase-field field directly encodes the material’s internal damage state as a scalar damage variable, whereas the electric potential field reflects the system’s electrostatic response, perturbed indirectly by the presence and evolution of cracks. This section provides a comparative evaluation of the two approaches, grounded in performance metrics, architectural sensitivity, interpretability, and practical implications for real-world applications.
@@
Tables 3 and 4 summarize the training and validation performance of the surrogate models, ranging from ResNet34 to ResNet152, comparing the phase-field-informed and electric potential-informed training strategies, respectively. As evident from these tables, models trained on electric potential data consistently outperformed their phase-field-based counterparts across all ResNet backbones. They not only achieved higher validation F1 and IoU scores but also demonstrated faster and more stable convergence. For instance, the ResNet34 model trained on electric potential data reached a validation IoU of 0.89 within just 50 epochs, whereas the ResNet34 model trained on phase-field data plateaued at an IoU of 0.76 after 75 epochs. This suggests that the electric potential fields provided richer, more learnable features for crack segmentation.
@@
A key insight from the comparative analysis is the role of network depth in segmentation performance across the two different training strategies. When trained on electric potential data, deeper ResNet variants–such as ResNet101 and ResNet152–show marginal but consistent improvements in validation accuracy compared to ResNet34. For instance, ResNet152 achieved a validation F1-score of 0.94 and a validation IoU of 0.90, slightly outperforming ResNet34, which achieved 0.93 and 0.89, respectively. These results indicate that deeper architectures can extract and utilize subtle spatial cues embedded in the smooth gradients of electric potential fields.
@@
In contrast, no such improvement is observed with deeper networks when trained on phase-field data. Despite the higher representational capacity of ResNet101 and ResNet152, their performance on phase-field data either stagnates or slightly declines relative to ResNet34. For example, ResNet34 outperformed all other variants with a validation F1-score of 0.82 and an IoU of 0.75, while deeper models failed to surpass this benchmark despite significantly longer training times. This suggests that the highly localized and discontinuous nature of phase-field crack images may not benefit from deeper architectures, which are generally more effective at learning hierarchical or global representations.
@@
The implication is that for phase-field-based segmentation tasks, increasing model depth yields diminishing or even negative returns, potentially due to overfitting to noise or sharp transitions in the data. On the other hand, electric potential data–with their smooth, continuous gradients–support better feature extraction in deeper networks, albeit with marginal gains.
Interpretability and physics-grounded representation
@@
While the phase-field approach is directly rooted in the mechanics of fracture, its effectiveness in training CNNs is often constrained by the highly localized and discontinuous nature of the damage variable. In contrast, the electric potential distribution represents a continuous, global field influenced by both material fracture and the underlying structural geometry. This yields significant advantages for deep learning; specifically, electric potential fields offer context-rich features by providing diffuse gradients that inherently reflect the presence of cracks, enabling the network to capture complex spatial dependencies more effectively.
The CNN’s ability to segment cracks from these potential maps is physically grounded in the ”insulating crack” boundary condition within the electromechanical framework. Regions where the damage variable $\mathfrak{d}\approx 1$ behave as low-permittivity channels. Physically, while the electric field penetrates the crack volume–modeled here as a medium with low but non-zero permittivity–the sharp contrast with the dielectric bulk matrix generates distinct field perturbations. Consequently, the electric potential and electric field are forced to bend sharply around the crack, creating high-gradient ”corridors” and distorted equipotential contours that serve as unique physical signatures of the crack path.
Consistent with the findings in Tan et al. [2023], the local field distribution is highly sensitive to the electrical boundary conditions at the crack faces. While the present study adopts a strong permittivity degradation with fixed electrode potentials, different physical regimes–such as short-circuit or partially conducting cracks–would yield distinct field signatures. Because these electro-mechanical distortions extend beyond the immediate crack tip, the training data provide greater spatial coverage and a more balanced class distribution, leading to enhanced numerical stability and faster convergence.
Ultimately, the network does not merely perform pattern recognition; it leverages these physically consistent field perturbations to reconstruct the high-fidelity ground-truth crack paths obtained from the phase-field simulations. The ability of the deep learning framework to generalize across varying electrical boundary conditions or to enforce perfectly insulating cracks remains a compelling direction for future work.
@@
Despite its superior performance, electric potential data-based learning is not without limitations. One challenge lies in interpretability: while phase-field values directly correlate with damage, interpreting voltage-induced field distortions as ”cracks” requires physical insight and a well-trained model. Furthermore, in complex geometries or low signal-to-noise regimes, voltage gradients may be too diffuse to pinpoint precise crack boundaries, potentially affecting precision.
A significant physical limitation arises in multi-crack scenarios due to the ”shielding effect” inherent in global elliptic fields like the electric potential. As observed in Figure 3, while the framework accurately captures dominant fracture paths, it may overlook secondary satellite cracks or non-propagating branches. Because the electric potential is a global field, its distribution is primarily dictated by the most significant geometric discontinuities. Dominant cracks can physically mask the signatures of smaller branches, leaving them with insufficient field gradients for the network to reliably segment. This suggests that while the current modality is highly sensitive to overall structural connectivity, it may require augmentation–such as multi-modal sensing incorporating local strain fields–to resolve fine-scale secondary cracking events.
@@
3.6 Computational efficiency and speed-up analysis
A primary motivation for developing the hybrid deep learning framework is to overcome the prohibitive computational cost associated with high-fidelity, fully coupled electromechanical phase-field simulations. To evaluate the efficiency of the proposed surrogate model, we conducted a direct comparison of the execution times for a single crack propagation case.
The reference high-fidelity simulations, performed using the finite element method on a workstation equipped with 32 CPU cores, required approximately 15 seconds per case to converge. In contrast, once trained, the ResNet-U-Net model completes a single inference (prediction) in approximately 0.1 seconds on an NVIDIA RTX 3090 GPU. This represents a significant 150-fold speed-up in prediction time.
Critically, this reduction in computational burden is achieved while maintaining high spatial accuracy, as evidenced by an IoU of approximately 0.9 for the electric potential-based predictions. Such a dramatic increase in efficiency demonstrates the potential of the proposed framework for real-time applications, such as structural health monitoring and iterative design optimization.